Operations Report – December 11th

Crew 184 Operations Report  12/11/2017

SOL: 10

Name of person filing report: HuntJ

Non-nominal systems:   EVA suits #5 / #6 (see below)

Notes on non-nominal systems: 

Generator (hours run): Generator turned off at 10:05 AM, nominal response.  Turned On at 5:45pm.

Solar— SOC @ 8:00am: 100%

              SOC @ 10:05am: 100%

             SOC @ 1:30 pm: 100%

             SOC @ 5:45pm:  65% (Generator Turned On)

Diesel –  99%

Propane –  Assume 69 percent volume (Due to trends from previous days)

Ethanol Free Gasoline (5 Gallon containers for ATV) –   1.5 Gallons

Water (trailer) –   400. gallons

Water (static) –   400 gallons

Trailer to Static Pump used –  Yes

Water (loft) – Static to Loft Pump used – Yes

Water Meter:  47 gallons

Toilet tank emptied: No

ATV’s Used: None,

Oil Added: No

ATV Fuel Used:  None

# Hours the ATVs were Used today: 00:00 Hours

Notes on ATVs: No ATVs used today due to no EVA.

Deimos rover used: No

Hours:

Beginning charge:

Ending charge:

Currently charging: Yes

Sojourner rover used:  ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: No

Hours:

Beginning charge:

Ending charge:

Currently charging: Yes

Opportunity rover used: No

Hours:

Beginning charge:

Ending charge:

Currently charging: Yes

Curiosity rover used: No

Hours:

Beginning charge:

Ending charge:

Currently charging: Yes

HabCar used and why, where?  NO

General notes and comments: Power is back and we’re back in sim!!

Summary of internet: 1.0 GB of Bandwidth remaining

Summary of suits and radios:

Radios: Nominal after getting charged post HAB powerdown.  Crew has primarily retired old model radios to preferentially use the new radios delivered by Mission Control.  Communication is much crisper / easier to understand on new radios, particularly when on EVA.

EVA Suits:

EVA 5: Fan does not turn on.

EVA 6: Fan does not turn on.  One of the two black straps holding the Helmet Collar in place to the suit is missing, resulting in helmet sitting incorrectly when in use and forcing the EVA crewmembers head down.

EVA 8: Right latch securing the back panel is loose. Was able to re secure by threading a  zip tie through the latch and tieing it down. After a quick vibration test (read: Shaking the suit a little) everything held steady. EVA 8 is operational at this time.

Summary of Hab operations:  Nominal

Summary of GreenHab operations: See Greenhab report

Summary of ScienceDome operations:  Currently being cleaned, Washvac assembled, Mission Control organizing Science Dome today

Summary of RAM operations: Not Operational

Summary of health and safety issues:  All Healthy

Questions, concerns and requests to Mission Support:

  1. Request more ATV Fuel. Can probably get through 2 more EVAs before exhausting remaining reserves.
  2. Burnable and Non burnable trash in airlock ready for pickup.